PROBLEM 2
Elder is paid a monthly salary of $2,250. Overtime is paid for hours
beyond 40 in each workweek. One week, Elder works 7 hours overtime.
Elder’s gross pay for the week is

PROBLEM 3
Ides receives 16 cents for every unit produced. Ides produces 2,976
pieces in a 43-hour workweek. For overtime, Ides is paid a sum equal to
one-half the regular hourly pay rate multiplied by the number of overtime
hours. Ides’ total piecework and overtime earnings are

PROBLEM 4
Kerr receives an annual $25,700 base salary for working the territory in
Arizona. A quota of $900,000 in sales has been set for that state. Kerr
receives an 8 percent commission on all sales in excess of $900,000.
This year, the sales are $965,000. The total earnings due Kerr this year
are

PROBLEM 5
Stacy Forvour is a salaried employee who works fluctuating workweeks.
She is paid $680 per workweek. This week, she worked 46 hours.
Forvour’s total gross pay if her employer uses the special half-rate
(based on total hours worked) for overtime pay is
Answer #1

Problem 2 Answer:

Elder's annual income = $2250*12 = $27000

Income per week = $27000/52 weeks = $519.23

Hourly Income = $519.23/40 = $ 12.98

Gross Pay for the week:

Salary = 519.23

Overtime = ($12.98*7 hours overtime * 1.5) = 136.29

Gross pay per week = $655.52

Problem 3 Answer:

Earnings for the production = 2,976 × $0.16 = $476.16

Hourly rate of Ides = $476.16/43 hours =$11.07

Computaion :

Piecework Earmings = $476.16

Overtime earnings= $11.07 × 0.5*3overtime hours = $16.62

Total of Piecework and ovetime earnings = $492.78

Problem 4 Answer:

The total earnings due Kerr this year:

Basic Annual salary = $25700

8% Commission on sales in excess of $900000 = ($965000-$900000)*0.08 = $5200

Total earnings = $25700+$5200= $30900

Problem 5 Answer:

Stacy Forvour earnings per workweek = $680

Earnings per hour = $680/46 hours = $14.78

Overtime earnings = ($14.78*0.5*6 overtime hours) = $44.34

Total Earnings per workweek = $680+$44.34 = $724.34